# A SAM-guided Two-stream Lightweight Model for Anomaly Detection

### v1

@@ Abstract.
In industrial anomaly detection, model efficiency and mobile-friendliness become the primary concerns in real-world applications. Simultaneously, the impressive generalization capabilities of Segment Anything (SAM) have garnered broad academic attention, making it an ideal choice for localizing unseen anomalies and diverse real-world patterns. In this paper, considering these two critical factors, we propose a SAM-guided Two-stream Lightweight Model for unsupervised anomaly detection (STLM) that not only aligns with the two practical application requirements but also harnesses the robust generalization capabilities of SAM. We employ two lightweight image encoders, i.e., our two-stream lightweight module, guided by SAM’s knowledge. To be specific, one stream is trained to generate discriminative and general feature representations in both normal and anomalous regions, while the other stream reconstructs the same images without anomalies, which effectively enhances the differentiation of two-stream representations when facing anomalous regions. Furthermore, we employ a shared mask decoder and a feature aggregation module to generate anomaly maps. Our experiments conducted on MVTec AD benchmark show that STLM, with about 16M parameters and achieving an inference time in 20ms, competes effectively with state-of-the-art methods in terms of performance, 98.26% on pixel-level AUC and 94.92% on PRO. We further experiment on more difficult datasets, e.g., VisA and DAGM, to demonstrate the effectiveness and generalizability of STLM. Codes are available online at https://github.com/StitchKoala/STLM.
@@ 1. Introduction
Image anomaly detection and localization tasks have attracted great attention in various domains, e.g., industrial quality control (Bergmann et al., 2019a; Roth et al., 2022), medical diagnoses (Fernando et al., 2021), and video surveillance (Liu and Ma, 2019; Guo et al., 2017). These tasks aim to discriminate both abnormal images and anomalous pixels in images according to previously seen normal or pseudo-anomalous samples during training. However, anomaly detection and localization are especially hard, as anomalous images occur rarely and anomalies can vary from subtle changes to large defects such as broken parts. Creating a dataset that includes sufficient anomalous samples with all possible anomaly types for training is a formidable challenge. Therefore, most methods have turned to tackling AD tasks with unsupervised models (Tien et al., 2023; Deng and Li, 2022; Chen et al., 2020), by only relying on normal samples, which are of great significance in practice.
@@
To achieve this, the utilization of a memory bank is proposed, in which a core set stores features extracted from the pre-trained backbone, to calculate a patch-level distance between the core set and the sample to detect anomalies (Yao et al., 2023a; Zhang et al., 2023d). However, these methods, which involve creating memory banks, come at the cost of increased computational complexity and large memory space.
Recent efforts have been directed towards proposing effective anomaly detection methods, with particular emphasis on reconstruction-based methods (Chen et al., 2023; Vasilev et al., 2020; Zhang et al., 2023b), which are favored for their simplicity and interpretability. These methods typically employ autoencoders (Bergmann et al., 2019b; Sarafijanovic-Djukic and Davis, 2019; Zhou and Paffenroth, 2017; Shi et al., 2021),variational autoencoders (Vasilev et al., 2020; Liu et al., 2020), or generative adversarial networks (Goodfellow et al., 2014). Based on the assumption that the reconstruction error of normal regions is lower than that of abnormal ones, they compare the test images with their reconstructed counterparts. However, recent studies have revealed that deep models can generalize so effectively that they accurately restore anomalous regions (Zavrtanik et al., 2021b), undermining the effectiveness of anomaly detection tasks.
To address this challenge, memory modules have been integrated into reconstruction-based methods (Park et al., 2020; Gong et al., 2019). These modules store representations of normal images, and the representations of test images are used as queries to retrieve the most relevant memory items for reconstruction. However, they suffer from high memory requirements and search times, making them impractical for real-world industrial applications.
The knowledge distillation (KD) (Hinton et al., 2015) based framework has shown its effectiveness in anomaly detection and localization (Tien et al., 2023; Roth et al., 2022; Salehi et al., 2021). For example, Salehi et al.  (Salehi et al., 2021) established an S-T network pair where knowledge is transferred from the teacher to the student. The underlying hypothesis is that the student network sees only the normal samples in the training stages, leading to generating out-of-distribution representations with anomalous queries during inference. However, the statement does not always hold true due to architectural similarities and shared data flow (Deng and Li, 2022). To overcome this limitation, DeSTSeg (Zhang et al., 2023c) introduces a denoising student network to generate distinct feature representations from those of the teacher when handling anomalous inputs.
@@
After processing the data, the TLM is introduced. First, instead of directly utilizing the pre-trained teacher network during inference in KD, we suggest distilling the comprehensive knowledge from the fixed SAM to a student stream as the lightweight “teacher”, called the plain student stream, which is more precise and generalized for representing features related to anomaly detection (AD) task. Second, inspired by (Zhang et al., 2023c), we incorporate the training of a denoising stream, also distilled from the SAM. The denoising student stream takes a pseudo-anomalous image as input, whereas the teacher SAM operates on the original, clean image. Our method effectively enhances the differentiation of two-stream features when dealing with anomalous regions. The impressive capabilities of SAM alleviate the generality problem typically associated with being opaque to genuine anomalies during training and various normal patterns in practical scenarios. In our method, we employ the encoder of the MobileSAM (Zhang et al., 2023a) as the backbone for our TLM’s encoder. Besides, we design a shared mask decoder and a feature aggregation module to generate anomaly maps.
@@ 2. Related Work
In this section, we review some related works to our work in the following part, including Deep Learning Methods for Anomaly Detection and Localization, Vision Foundation Models and Data augmentation strategies for AD tasks.
@@ 2.1. Deep Learning Methods for Anomaly Detection and Localization
Prior works of anomaly detection and localization have explored various methodologies. The method of image reconstruction is widely adopted, which posits that accurately reconstructing anomalous regions can be challenging due to their absence in the training samples. Generative models such as autoencoders (Bergmann et al., 2019b; Li et al., 2021a), variational autoencoders (Vasilev et al., 2020; Xu et al., 2023), and generative adversarial networks (Goodfellow et al., 2014; Peng and Qi, 2019) are also utilized to reconstruct normal images from anomalous ones (Akrami et al., 2022). Nonetheless, these methods face certain limitations, especially when reconstructing complex textures and patterns. Later methods use deep models to enhance the quality of reconstructing images (Zavrtanik et al., 2021a, b; Zhang et al., 2023b).
Recently, the utilization of a memory bank, in which a core set stores features extracted from the pre-trained backbone, calculates a patch-level distance between the core set and the sample to detect anomalies (Yao et al., 2023a). The studies in (Zhang et al., 2023d) reveal a method that focuses on learning feature residuals of varying scales and sizes between anomalous and normal patterns, which prove beneficial in accurately reconstructing the segmentation maps of anomalous regions. However, these methods, which involve creating memory banks, come at the cost of increased computational complexity.
@@ 2.2. Vision Foundation Models
Vision Foundation Models showcase remarkable zero-shot capabilities to address numerous downstream tasks. CLIP (Radford et al., 2021) can represent visual content using textual prompts and is trained on 400 million image-text pairs, which rivals the image classification accuracy of fully supervised works. Recently, SAM (Kirillov et al., 2023), which is trained on millions of annotated images, generates high-quality segmentation results for previously unseen images. However, foundation models often exhibit limitations in specific domains, including medical imaging and the field of anomaly detection (Ji et al., 2023). Since they are mostly transformer-based frameworks, they also have issues with computation complexity and inference latency at inference time (Pope et al., 2023). And they are mostly trained in natural images, they tend to prioritize foreground objects and may struggle to accurately segment small or irregular objects. Therefore, our work seeks to explore how to effectively harness the extensive knowledge provided by these off-the-shelf models to detect anomalies.
@@
2.3. Data augmentation strategies
Data augmentation strategies play a pivotal role in the field. One widely used technique involves simulating pseudo-anomalous data by introducing artificial anomalies into the provided anomaly-free samples. This method effectively transforms the one-class classification anomaly detection task into a supervised learning framework (Li et al., 2021b; Zavrtanik et al., 2021b; Collin and De Vleeschouwer, 2021). Classical anomaly simulation strategies, such as rotation (Gidaris et al., 2018) and cutout (DeVries and Taylor, 2017), have exhibited limitations in effectively identifying subtle anomalous patterns. A simple yet effective strategy called CutPaste (Li et al., 2021b) generates pseudo-anomalies by copying and pasting a rectangular region from one part of the image to another. Since the model focuses on detecting local features such as edge discontinuity and texture perturbations, it may fall short when dealing with detecting and localizing larger defects and global structural anomalies. Also,it typically generate only one type of pseudo-anomaly, limiting their adaptability to various anomaly types commonly encountered in industrial settings.
More recent researchers have adopted the innovative ideas presented in (Zavrtanik et al., 2021a) for anomaly simulation, and we have also integrated pseudo-anomalous data with the capacity to generalize across diverse types of unseen defects. However, the introduction of pseudo-anomalies onto normal images renders these models more sensitive to anomalous regions, factors like noisy backgrounds or fluctuating illumination in test images may lead to distractions and increase false positive detections (Collin and De Vleeschouwer, 2021). It is noteworthy that, SimpleNet counterfeits anomaly features by introducing Gaussian noise to normal features, a process that requires careful tuning of appropriate hyperparameters.
@@ 3. Method
Together with the pseudo anomalies introduced into normal training images with predefined probabilities (in Section 3.1) to maintain balanced data distribution, we propose the SAM-guided Two-stream Lightweight Model for unsupervised anomaly detection (STLM) to effectively generate an anomaly map for anomaly detection and localization. As illustrated in Figure 2, the two-stream lightweight model (TLM, in Section 3.2), consisting of a plain student stream and a denoising student stream, distills different information from a fixed SAM (Kirillov et al., 2023) during training. Consequently, a feature aggregation module (FA, in Section 3.3) is trained to fuse anomaly features with the aid of additional supervision signals. For inference (in Section 3.4), anomaly maps are generated solely using the TLM and the FA module. In the subsequent sections, we will provide a detailed description.
@@ 3.1. Pseudo Anomaly Generation
where $N$ is the normal sample, $A$ is the external data source an arbitrary image from, $\bar{M}$ is the inverse of $M$, $\odot$ means the element-wise multiplication operation. $\beta$ is the opacity parameter for a better combination of abnormal and normal regions. Figure 3 shows the pseudo-anomalous images generated by this strategy. In practice, we use a probability of 0.5 to decide if this generation method is activated or not, with the ablation study shown in Figure 10. Notably, the generation process is conducted in real-time training.
Figure 3. Examples of pseudo-anomalous images generated
@@ 3.2. Two-Stream Lightweight Model
In traditional KD, the teacher network typically employs a large and deep expert network with extensive capabilities, while the student network adopts a similar neural network structure to that of the teacher (Bergmann et al., 2019b). Although this method enhances precision compared to early works, anomaly detection does not always hold true due to architectural similarities and shared data flow. Despite (Zhang et al., 2023c) introducing a denoising student network to address this issue, it still fails to consider that the teacher network struggles to accurately represent unseen normal textures and patterns with some variations compared to the training set. And the denoising student network has generality limitations of being opaque to genuine anomalies in the training stage.
@@ 3.2.1. Mobile Distillation
First, we follow the idea of encouraging the student network to generate anomaly-specific features that diverge from those of the teacher network adopted by (Zhang et al., 2023d, c). This strategy effectively helps overcome the limitations posed by network architecture similarities and identical data flows. However, as (Zhang et al., 2023b) said, the test data can have large distribution shifts arising in many real-world applications and may correspond to multiple normal patterns, leading to imprecise representations of the teacher network. The plain student stream is trained to produce discriminative and general feature representations in both normal and anomalous regions as a superior alternative to the teacher network, demonstrated in Section 4.4. Similarly guided by SAM’s knowledge, the denoising student stream aligns its feature representations with those of the same images without any corruption, which effectively enhances the differentiation of the representations from two student streams when dealing with anomalous regions.
@@
Figure 4. Comparison of different architectures between (a) the original two-separated-decoders version and (b) one-shared-decoder version. We observe that the one-shared-decoder version, while economical in terms of parameters and computation time, makes a satisfactory result with only a slight reduction compared to employing two separated decoders.
Consequently, SAM serves as the teacher network for the TLM. The output feature maps are extracted from the two-layer decoder of SAM, shown in Figure 4 (a). With no need for segmentation outcomes, we only use the Two-Way-Transformer blocks of the mask decoder, and introduce our FA module. The output features from the pre-trained SAM with normal images as input are denoted as $T_{P}^{1}$ and $T_{P}^{2}$, respectively. Similarly, those originating from pseudo-anomalous images are labeled $T_{D}^{1}$ and $T_{D}^{2}$. TLM adopts the image encoder of mobileSAM (Zhang et al., 2023a), i.e., ViT-Tiny (Wu et al., 2022), and their output features are denoted as $S_{P}^{1}$, $S_{P}^{2}$, $S_{D}^{1}$, and $S_{D}^{2}$, respectively.
@@
Figure 5. Comparison between one shared decoder and two separated decoders.

### v2

@@ Abstract.
In industrial anomaly detection, model efficiency and mobile-friendliness become the primary concerns in real-world applications. Simultaneously, the impressive generalization capabilities of Segment Anything (SAM) have garnered broad academic attention, making it an ideal choice for localizing unseen anomalies and diverse real-world patterns. In this paper, considering these two critical factors, we propose a SAM-guided Two-stream Lightweight Model for unsupervised anomaly detection (STLM) that not only aligns with the two practical application requirements but also harnesses the robust generalization capabilities of SAM. We employ two lightweight image encoders, i.e., our two-stream lightweight module, guided by SAM’s knowledge. To be specific, one stream is trained to generate discriminative and general feature representations in both normal and anomalous regions, while the other stream reconstructs the same images without anomalies, which effectively enhances the differentiation of two-stream representations when facing anomalous regions. Furthermore, we employ a shared mask decoder and a feature aggregation module to generate anomaly maps. Our experiments conducted on MVTec AD benchmark show that STLM, with about 16M parameters and achieving an inference time in 20ms, competes effectively with state-of-the-art methods in terms of performance, 98.26% on pixel-level AUC and 94.92% on PRO. We further experiment on more difficult datasets, e.g., VisA and DAGM, to demonstrate the effectiveness and generalizability of STLM. Codes are available online at https://github.com/Qi5Lei/STLM.
Vision Foundation Models,Lightweight model,Unsupervised learning
@@
The work is supported by NSFC Program (Grants No. 62206052, 62125602, 62076063), China Postdoctoral Science Foundation (Grants No. 2024M750424), Supported by the Postdoctoral Fellowship Program of CPSF (Grant No. GZC20240252), and Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2024ZB242).
copyright: acmlicensedjournalyear: 2024submissionid: 2024-0694ccs: Computing methodologies Anomaly detectionccs: Computing methodologies Neural networksccs: Computing methodologies Object detection
@@ 1. Introduction
Image anomaly detection and localization tasks have attracted great attention in various domains, e.g., industrial quality control (Bergmann et al., 2019a; Roth et al., 2022), medical diagnoses (Fernando et al., 2021; SHI and CHE, 2023), and video surveillance (Liu and Ma, 2019; Guo et al., 2017). These tasks aim to discriminate both abnormal images and anomalous pixels in images according to previously seen normal or pseudo-anomalous samples during training. However, anomaly detection and localization are especially hard, as anomalous images occur rarely and anomalies can vary from subtle changes to large defects such as broken parts. Creating a dataset that includes sufficient anomalous samples with all possible anomaly types for training is a formidable challenge. Therefore, most methods have turned to tackling AD tasks with unsupervised models (Tien et al., 2023; Deng and Li, 2022; Chen et al., 2020), by only relying on normal samples, which are of great significance in practice.
@@
To achieve this, the utilization of a memory bank is proposed, in which a core set stores features extracted from the pre-trained backbone, to calculate a patch-level distance between the core set and the sample (Yao et al., 2023a; Zhang et al., 2023c). However, these methods, which involve creating memory banks, come at the cost of increased computational complexity and large memory space.
Recent efforts have been directed towards proposing effective anomaly detection methods, with particular emphasis on reconstruction-based methods (Chen et al., 2023), which are favored for their simplicity and interpretability. These methods typically employ autoencoders (Bergmann et al., 2019b; ZHU et al., 2023; Zhou and Paffenroth, 2017),variational autoencoders (Liu et al., 2020), or generative adversarial networks (Goodfellow et al., 2014). Based on the assumption that the reconstruction error of normal regions is lower than that of abnormal ones, they compare the test images with their reconstructed counterparts. However, recent studies have revealed that deep models can generalize so effectively that they accurately restore anomalous regions (Zavrtanik et al., 2021b), undermining the effectiveness of anomaly detection tasks.
To address this challenge, memory modules have been integrated into reconstruction-based methods (Gong et al., 2019). These modules store representations of normal images, and the representations of test images are used as queries to retrieve the most relevant memory items for reconstruction. However, they suffer from high memory requirements and search times.
The knowledge distillation (KD) (Hinton et al., 2015) based framework has shown its effectiveness in anomaly detection and localization (Tien et al., 2023; Roth et al., 2022; Salehi et al., 2021). For example, Salehi et al.  (Salehi et al., 2021) established an S-T network pair where knowledge is transferred from the teacher to the student. The underlying hypothesis is that the student network sees only the normal samples in the training stages, leading to generating out-of-distribution representations with anomalous queries during inference. However, the statement does not always hold true due to architectural similarities and shared data flow (Deng and Li, 2022). To overcome this limitation, DeSTSeg (Zhang et al., 2023b) introduces a denoising student network to generate distinct feature representations from those of the teacher when handling anomalous inputs.
@@
After processing the data, the TLM is introduced. First, instead of directly utilizing the pre-trained teacher network during inference in KD, we suggest distilling the comprehensive knowledge from the fixed SAM to a student stream as the lightweight “teacher”, called the plain student stream, which is more precise and generalized for representing features related to anomaly detection (AD) task. Second, inspired by (Zhang et al., 2023b), we incorporate the training of a denoising stream, also distilled from the SAM. The denoising student stream takes a pseudo-anomalous image as input, whereas the teacher SAM operates on the original, clean image. Our method effectively enhances the differentiation of two-stream features when dealing with anomalous regions. The impressive capabilities of SAM alleviate the generality problem typically associated with being opaque to genuine anomalies during training and various normal patterns in practical scenarios. In our method, we employ the encoder of the MobileSAM (Zhang et al., 2023a) as the backbone for our TLM’s encoder. Besides, we design a shared mask decoder and a feature aggregation module to generate anomaly maps.
@@ 2. Related Work
In this section, we review some related works to our work, including Deep Learning Methods for Anomaly Detection and Localization, Vision Foundation Models and Lightweight Models.
@@ 2.1. Deep Learning Methods for Anomaly Detection and Localization
Prior works of anomaly detection and localization have explored various methodologies. The method of image reconstruction is widely adopted, which posits that accurately reconstructing anomalous regions can be challenging due to their absence in the training samples. Generative models such as autoencoders (Bergmann et al., 2019b; Li et al., 2021a), variational autoencoders (Xu et al., 2023), and generative adversarial networks (Goodfellow et al., 2014; Peng and Qi, 2019) are also utilized to reconstruct normal images from anomalous ones. Nonetheless, these methods face certain limitations, especially when reconstructing complex textures and patterns. Later methods use deep models to enhance the quality of reconstructing images (Zavrtanik et al., 2021a, b).
Recently, the utilization of a memory bank, in which a core set stores features extracted from the pre-trained backbone, calculates a patch-level distance between the core set and the sample to detect anomalies (Yao et al., 2023a). The studies in (Zhang et al., 2023c) reveal a method that focuses on learning feature residuals of varying scales and sizes between anomalous and normal patterns, which prove beneficial in accurately reconstructing the segmentation maps of anomalous regions. However, these methods, which involve creating memory banks, come at the cost of increased computational complexity.
@@ 2.2. Vision Foundation Models
Foundation models (Bommasani et al., 2021; Brown et al., 2020) initially emerged, that large language models such as the GPT series (Li et al., 2023b) showcased impressive zero-shot generalization (Lin et al., 2022; Zheng et al., 2022; GAO et al., 2023) to unseen tasks and data. This led to the development of prompt-based learning works (Houlsby et al., 2019; Hu et al., 2022) aimed at helping these pre-trained models generalize to downstream tasks instead of fine-tuning their internal parameters (Hinterstoisser et al., 2018) for better transfer learning.
For vision-based foundation models (Ji et al., 2023; Wang et al., 2023; Zou et al., 2024), prompt engineering (Zhou et al., 2023)— freezing the pre-trained model— was first explored in vision-language models like CLIP. CLIP (Radford et al., 2021), which was trained on 400 million image-text pairs, can represent visual content through textual prompts. It was demonstrated strong zero-shot performance on major benchmark datasets. However, CLIP’s performance diminishes when dealing with images not included in its pre-training datasets, highlighting the impact of training data size and distribution on downstream task performance. Although CLIP and CLIP-style models facilitate vision tasks by linking images with descriptive text to capture scene-level semantics, they struggle with identifying instance-level semantics within scenes.
DINO (Caron et al., 2021) utilizes a self-supervised student-teacher framework based on vision transformers (Dosovitskiy et al., 2021). It learns robust and high-quality features for downstream tasks due to employing a multi-view strategy. Recently, further advances have been shown by SAM (Kirillov et al., 2023) and DINOv2 (Oquab et al., 2024). DINOv2 integrates ideas from DINO with patch-level reconstruction methods (Zhou et al., 2022). SAM, trained on millions of annotated images, generated high-quality segmentation results for previously unseen images. Their capabilities have been evaluated across numerous tasks, showcasing impressive zero-shot performance that is often competitive with or even superior to prior fully supervised methods. Additionally, we aim to leverage SAM’s exceptional segmentation strength for our AD tasks.
However, foundation models often exhibit limitations in specific domains, including medical imaging and anomaly detection (Ji et al., 2023). Since they are mostly transformer-based frameworks, they also have issues with computation complexity and inference latency at inference time (Pope et al., 2023). And they are mostly trained in natural images, they tend to prioritize foreground objects and may struggle to accurately segment small or irregular objects. Therefore, our work seeks to explore how to effectively harness the extensive knowledge these off-the-shelf models provide.
2.3. Lightweight Models
While deep learning offers significant advantages for feature extraction, traditional neural networks often demand substantial computational power and memory resources during training. In recent years, there has been notable progress in developing lightweight models, particularly for deployment on resource-constrained devices. To address these challenges, techniques like model pruning (Wang et al., 2024; Liu et al., 2019), quantization (Yao et al., 2021; Liu et al., 2021), and knowledge distillation (Hinton et al., 2015; Salehi et al., 2021) have been explored to reduce model size and computational complexity while maintaining performance. CNN-based lightweight architectures such as MobileNet (Howard et al., 2017), ShuffleNet (Zhang et al., 2018), and CSPNeXt (Chen et al., 2024) have introduced innovations like depthwise separable convolutions and group convolutions, significantly reducing inference time and memory requirements.
Despite these advancements, challenges persist particularly in unsupervised tasks like anomaly detection, where accuracy, efficiency and generalization are crucial. State-of-the-art CNN-based methods, such as SimpleNet (Liu et al., 2023) and DeSTSeg, achieve satisfactory performance but fall short in generalization across multiple datasets. Transformers, with their large amount of parameters, offer high interpretability through self-attention mechanisms (Cao et al., 2024). It also captures global dependencies, in contrast to CNNs which tend to emphasize local features. These properties help Transformers to a more suitable architecture for unsupervised tasks, particularly in training generalized vision models on large-scale datasets, such as SAM and DINOv2. Recent works, proposing efficient Transformers, have further enhanced their applicability in resource-constrained environments.
@@ 3. Method
Together with the pseudo anomalies introduced into normal training images with predefined probabilities (in Section 3.1) to maintain balanced data distribution, we propose the SAM-guided Two-stream Lightweight Model for unsupervised anomaly detection (STLM) to effectively generate an anomaly map for anomaly detection and localization. As illustrated in Figure 2, the two-stream lightweight model (TLM, in Section 3.2), consisting of a plain student stream and a denoising student stream, distills different information from a fixed SAM (Kirillov et al., 2023) during training. Consequently, a feature aggregation module (FA, in Section 3.3) is trained to fuse anomaly features with the aid of additional supervision signals. For inference (in Section 3.4), anomaly maps are generated solely using the TLM and the FA module.
@@ 3.1. Pseudo Anomaly Generation
where $N$ is the normal sample, $A$ is the external data source an arbitrary image from, $\bar{M}$ is the inverse of $M$, $\odot$ means the element-wise multiplication operation. $\beta$ is the opacity parameter for a better combination of abnormal and normal regions. In practice, we use a probability of 0.5 to decide if this generation method is activated or not, with the ablation study shown in Figure 7(c). Notably, the generation process is conducted in real-time training.
@@ 3.2. Two-Stream Lightweight Model
In traditional KD, the teacher network typically employs a large and deep expert network with extensive capabilities, while the student network adopts a similar neural network structure to that of the teacher (Bergmann et al., 2019b), which cannot learn about anomalous regions, leading to discrepancies with the teacher network during testing. However, extensive experiments have shown that this structure can cause collapsing (constantly generating the same normal image). Therefore, the reconstructed image is used for anomaly discrimination, while the teacher generator acts as a regularizer to prevent this issue. Reconstructing anomalous regions as normal reduces the student network’s ability to “discriminate” anomalous areas, thereby increasing the recognition difference between the teacher and student networks. Despite (Zhang et al., 2023b) introducing a denoising student network to address this issue, it still fails to consider that the teacher network struggles to accurately represent unseen normal textures and patterns with some variations compared to the training set. The denoising student network has generality limitations of being opaque to genuine anomalies while training.
@@ 3.2.1. Mobile Distillation
First, the TLM is a variation of the commonly used Student-Teacher(ST) distillation network in Anomaly Detection. In a conventional ST network, the student network, trained only on normal images, cannot learn about anomalous regions, leading to discrepancies with the teacher network during testing. However, extensive experiments have shown that this structure can cause collapse (constantly generating the same normal image). Therefore, the reconstructed image is used for anomaly discrimination, while the teacher generator acts as a regularizer to prevent this issue. Reconstructing anomalous regions as normal reduces the student network’s ability to “discriminate” anomalous areas, thereby increasing the recognition difference between the teacher and student networks. This approach has also been studied in (Bergmann et al., 2020; Zhang et al., 2023b). Specifically, the plain student stream is trained to produce discriminative and general feature representations in both normal and anomalous regions as a superior alternative to the teacher network, demonstrated in Section 4.4. Similarly guided by SAM’s knowledge, the denoising student stream aligns its feature representations with those of the same images without any corruption, which effectively enhances the differentiation of the representations from two student streams when dealing with anomalous regions.
Notably, based on the results from Section 4.4, where one-stage training outperforms two-stage training, we explain why it performs better: the FA and TLM modules being trained simultaneously. On the one hand, the trained plain stream, compared to the frozen SAM teacher network, exhibits greater adaptability to other parts of TLM, which generates better intermediate results feeding to the FA module. On the other hand, segmentation knowledge related to anomaly detection images is feedback through the gradients of the FA module, back into both streams to better recognize normal regions. Only the denoising student is improved without distillation training for SAM, resulting in differences in normal regions and affecting the experimental results.
@@
Consequently, SAM serves as the teacher network for the TLM. The output feature maps are extracted from the two-layer decoder of SAM, shown in Figure 2 (c). With no need for segmentation outcomes, we only use the Two-Way-Transformer blocks of the mask decoder, and introduce our FA module. The output features from the pre-trained SAM with normal images as input are denoted as $T_{D}^{1}$ and $T_{D}^{2}$, respectively. Similarly, those originating from pseudo-anomalous images are labeled $T_{P}^{1}$ and $T_{P}^{2}$. TLM adopts the image encoder of mobileSAM (Zhang et al., 2023a), i.e., ViT-Tiny (Wu et al., 2022), and their output features are denoted as $S_{P}^{1}$, $S_{P}^{2}$, $S_{D}^{1}$, and $S_{D}^{2}$, respectively.